According to Masterfoods, the company that manufactures M&M's, 12% of peanut M&M's are brown, 15% are yellow, 12% are red, 23% are blue, 23% are orange and 15% are green.
Round to 4 decimal places
a. Compute the probability that a randomly selected peanut M&M is not yellow.
b. A randomly selected peanut M&M is green or brown.
c. That two randomly selected peanut M&M's are both brown.
d. If you randomly select three peanut M&M's, what is the probability that none are brown.
e. If you randomly select three peanut M&M's, what is the probability that at least one of them is brown.

Answer by ikleyn(53937)

.

(a)  add everything except yellow.



(b)  add green and brown.



(c)  0.12 * 0.12 = 0.0144 = 1.44%.



(d)  (1-0.12)*(1-0.12)*(1-0.12) = use your calculator.



(e)  1 - (1-0.12)*(1-0.12)*(1-0.12) = use your calculator.

Solved and explained.